# Loading non-Maxwellian Velocity Distributions in Particle Simulations

### v1

@@ VI Ring and Shell distributions
$\mathrm{CDF}_{\mathrm{shell}}(v)$ $=1-\frac{\frac{(v+V)}{\theta}\exp\left(-\frac{(v-V)^{2}}{\theta^{2}}\right)+\sqrt{\pi}\left(\frac{V^{2}}{\theta^{2}}+\frac{1}{2}\right)\mathrm{erfc}{\left(\frac{v-V}{\theta}\right)}}{A_{3}(\frac{V}{\theta})}$ (82)
@@ VII Ring and Shell Maxwellians
where $V$ is the ring velocity and $I_{0}(x)$ is the zeroth-order modified Bessel function of the first kind. We call this distribution the ring Maxwell distribution or the ring Maxwellian [^1] The original authors call the distribution the “pseudo Maxwellian,” because they study near-Maxwellian cases, $V\lesssim\theta_{\perp}$. In this study, we call it the ring Maxwellian to emphasize its role as a ring-type distribution.. Note that we employ a different convention for the characteristic velocity from the original paper, by a factor of $\sqrt{2}$. The phase-space density is concave at $v_{\perp}\simeq 0$ when $V\leq\theta_{\perp}$, while it has a ring-shaped structure when $\theta_{\perp}<V$ [42]. The pressures and energy density of the ring Maxwellian are as follows.
$P_{\parallel}=\frac{1}{2}N_{0}m\theta_{\parallel}^{2},~~P_{\perp}=\frac{1}{2}N_{0}m(V^{2}+\theta_{\perp}^{2}),~~\mathcal{E}=\frac{1}{4}N_{0}m(2V^{2}+\theta_{\parallel}^{2}+2\theta_{\perp}^{2}).$ (85)
In this study, we propose its shell-type extension, called the shell Maxwell distribution or the shell Maxwellian. As will be shown, the shell Maxwellian is spherically scattered from a seed Maxwellian, while the ring Maxwellian is axisymmetrically scattered. Its phase-space density, pressure, and energy density are given by
@@
This immediately gives Eq. (86). Maclaurin expansion of Eq. (86) gives further insights:
@@
This indicates that the shell Maxwellian has a peak at the center, and it no longer looks like a shell when $V\leq\sqrt{{3}/{2}}\theta\simeq 1.22\theta$. The phase-space density has a shell-shaped structure when $V>\sqrt{{3}/{2}}\theta$.
@@
$\dfrac{f_{\mathrm{ring}}(v_{\perp}){\,2\pi v_{\perp}}}{f_{\mathrm{rM}}(v_{\perp}){\,2\pi v_{\perp}}}$ $=\frac{1}{A_{2}(\frac{V}{\theta_{\perp}})}\dfrac{\exp\left(\frac{2v_{\perp}V}{\theta_{\perp}^{2}}\right)}{I_{0}\left(\frac{2v_{\perp}V}{\theta_{\perp}^{2}}\right)},~~~~~\dfrac{f_{\mathrm{shell}}(r){\,4\pi r^{2}}}{f_{\mathrm{sM}}(r){\,4\pi r^{2}}}=\frac{2rV\sqrt{\pi}}{\theta^{2}A_{3}(\frac{V}{\theta})}\left\{1-\exp\left(-\frac{4rV}{\theta^{2}}\right)\right\}^{-1},$ (96)
As evident in Figure 4(e), the two ring distributions are very similar around the peaks for $V=5\theta_{\perp}$. The phase-space densities are maximum at $v_{\perp}=5\theta_{\perp}$ for the ring distribution, and at $v_{\perp}\approx 4.95\theta_{\perp}$ for the ring Maxwellian. Also, in Figure 4(f), the two shell distributions are similar around the peaks for $V=5\theta$. The maximums are located at $v=5\theta$ for the shell distribution, and at $v\approx 4.9\theta$ for the shell Maxwellian. These results suggest that the ring and shell Maxwellians are good alternatives for the ring and shell distributions with Gaussian width. On the other hand, when $V\lesssim\theta$, the ring (shell) distribution and the ring (shell) Maxwellian are very different, as shown by the dotted lines. As already stated, the artificial boundary of the ring and shell distributions becomes problematic.
@@ VIII Other isotropic distributions
We present numerical procedures for two isotropic distributions. First, we discuss the super-Gaussian distribution, also known as the self-similar distribution. The distribution has long been discussed in laser plasmas [11, 26] but also useful in space plasmas [45]. It is given by
@@
Finally we consider the filled-shell distribution, which plays a role in a distant heliosphere. The phase-space density follows a power-law $\propto v^{p}$ inside a spherical shell ($v\leq V$),
@@
The spectral index varies from problem to problem. The so-called Vasyliunas–Siscoe model [44] employs p=3/2p=-3/2 [e.g., 52], but we find other choices in the literature [e.g., $p=1$; 14].
@@
Figure 5: The phase-space densities (black lines) and Monte Carlo results of $10^{6}$ particles (blue histograms) for (a) the super Gaussian distribution ($p=3$, $\theta=1$) and (b) the filled-shell distribution (p=3/2p=-3/2, $V=2$) are presented.
IX Discussion and Summary
@@
In this paper, we have provided numerical recipes for nine different VDFs: the $(r,q)$ distribution, the regularized kappa distribution, the subtracted kappa distribution, the ring distribution, the shell distribution, the ring Maxwellian, the shell Maxwellian, the super-Gaussian, and the filled-shell distributions. The recipes require only three elemental variates: the uniform variate, the normal variate, and the gamma variate. The algorithms are fully prescribed in tables, making it easy for users to implement them in the code.
@@
In Section III, we have discussed the $(r,q)$ distribution that generalizes the kappa and flattop distributions. One can use either of the beta-prime method and the piecewise rejection method, if one is aware of limitations of each. Sections IV and V have discussed the two advanced kappa distributions. For the regularized kappa distribution, the two rejection methods and their preferred parameter spaces are described. The simple procedure was proposed for a very recent model of the subtracted kappa distribution. The loss-cone filling parameter $\Delta$ is considered in Eq. (66). A previous procedure for the subtracted Maxwellian [54, Section 2] can be similarly modified to incorporate the filling parameter.
@@
In Section VII, we have introduced the ring Maxwellian and the shell Maxwellian. We would propose them as alternatives for the ring and shell distributions in Section VI. When the ring/shell velocity is larger than the thermal width, $V\gg\theta$, they are similar to the conventional ring and shell distributions. When $V\lesssim\theta$, the conventional distributions become somewhat artificial near $v=0$, and probably it does not make sense to employ them. In contrast, The (ring and shell) Maxwellians have no such problem. The new Maxwellians have other nice features. As presented, it is much easier to estimate pressures and energy density for the new Maxwellians than those for the conventional distributions. Also, Monte Carlo procedures for the new Maxwellians are much simpler than those for the conventional distributions, as shown in Tables 4 and 5. All these facts suggest that the ring or shell Maxwellians are favorable for modeling of heliospheric kinetic processes in ring or shell-shaped VDFs.
@@
For completeness, we have discussed two isotropic VDFs in Section VIII. Together with all the VDFs in this paper and in our earlier papers [53, 54], it is now possible to generate a wide variety of single-component VDFs in heliophysics for particle simulations. In advanced settings, VDFs may have multiple components, some of which are not approximated by well-known VDFs. For example, solar-wind electrons often contain core, halo, and strahl populations. In such cases, it may be necessary not only to prepare separate procedures for each component, but also to use a fully numerical inversion method [4]. This needs to be considered on a case-by-case basis.
@@
In summary, we have provided simple and efficient procedures for generating various VDFs in heliophysics research. We hope that this work will make kinetic modeling of heliospheric plasma processes easier than before.
@@ Appendix A Piecewise rejection method for log-concave distributions
In practice, the normalization constant can be dropped from $g(x)$, because subsequent discussion does not rely on it. The distribution has a maximum at the most probable speed:
@@
are the two switching points. As an example, the envelope functions are drawn by the dashed orange lines in Figures 4(a,b). The ratio of the areas below the envelopes is $\lambda(v_{L}):(x_{R}-x_{L}):\lambda(v_{R})$. Then a rejection algorithm can be constructed, as shown in Table 7. Surprisingly, the acceptance efficiency is $\gtrsim 86\%$ for $V\gtrsim\theta$ and $88\%$$89\%$ for $V\gtrsim 2\theta$ for both ring and shell distributions.
@@
$p_{L}\leftarrow\dfrac{\lambda_{L}}{S}$$p_{R}\leftarrow\dfrac{\lambda_{R}}{S}$$p_{C}\leftarrow 1-q_{L}-q_{R}$

### v2

@@ VI Ring and Shell distributions
$\mathrm{CDF}_{\mathrm{shell}}(v)$ $=1-\frac{\frac{(v+V)}{\theta}\exp\left(-\frac{(v-V)^{2}}{\theta^{2}}\right)+\sqrt{\pi}\left(\frac{V^{2}}{\theta^{2}}+\frac{1}{2}\right)\mathrm{erfc}{\left(\frac{v-V}{\theta}\right)}}{C_{3}(\frac{V}{\theta})}$ (85)
@@ VII Ring and Shell Maxwellians
where $V$ is the ring velocity and $I_{0}(x)$ is the zeroth-order modified Bessel function of the first kind. We call this distribution the ring Maxwell distribution or the ring Maxwellian [^1] The original authors call the distribution the “pseudo Maxwellian,” because they study near-Maxwellian cases, $V\lesssim\theta_{\perp}$. In this study, we call it the ring Maxwellian to emphasize its role as a ring-type distribution.. Note that we employ a different convention for the characteristic velocity from the original paper, by a factor of $\sqrt{2}$. The phase space density is concave at $v_{\perp}\simeq 0$ when $V\leq\theta_{\perp}$, while it has a ring-shaped structure when $\theta_{\perp}<V$ [45]. The pressures, the energy density, and the pressure anisotropy of the ring Maxwellian are as follows.
$P_{\parallel}=\frac{1}{2}N_{0}m\theta_{\parallel}^{2},~~P_{\perp}=\frac{1}{2}N_{0}m(V^{2}+\theta_{\perp}^{2}),$ (88)
$\mathcal{E}=\frac{1}{4}N_{0}m(2V^{2}+\theta_{\parallel}^{2}+2\theta_{\perp}^{2}),~~A=\frac{V^{2}+\theta_{\perp}^{2}}{\theta_{\parallel}^{2}}-1.$ (89)
In this study, we propose its shell-type extension, called the shell Maxwell distribution or the shell Maxwellian. As will be shown, the shell Maxwellian is spherically scattered from a seed Maxwellian, while the ring Maxwellian is axisymmetrically scattered. Its phase space density, pressure, and energy density are given by
@@
This immediately gives Eq. (90). The energy distribution resembles that of a drifting Maxwellian, discussed by Karlický et al. [16, Section 2]. Maclaurin expansion of Eq. (90) gives further insights:
@@
This indicates that the shell Maxwellian has a peak at the center, and it no longer looks like a shell when $V\leq\sqrt{{3}/{2}}\theta\simeq 1.22\theta$. The phase space density has a shell-shaped structure when $V>\sqrt{{3}/{2}}\theta$.
@@
$\dfrac{f_{\mathrm{ring}}(v_{\perp}){\,2\pi v_{\perp}}}{f_{\mathrm{rM}}(v_{\perp}){\,2\pi v_{\perp}}}$ $=\frac{1}{C_{2}(\frac{V}{\theta_{\perp}})}\dfrac{\exp\left(\frac{2v_{\perp}V}{\theta_{\perp}^{2}}\right)}{I_{0}\left(\frac{2v_{\perp}V}{\theta_{\perp}^{2}}\right)},~~~~~\dfrac{f_{\mathrm{shell}}(r){\,4\pi r^{2}}}{f_{\mathrm{sM}}(r){\,4\pi r^{2}}}=\frac{2rV\sqrt{\pi}}{\theta^{2}C_{3}(\frac{V}{\theta})}\left\{1-\exp\left(-\frac{4rV}{\theta^{2}}\right)\right\}^{-1},$ (100)
As evident in Figure 4(e), the two ring distributions are very similar around the peaks for $V=5\theta_{\perp}$. The phase space densities are maximum at $v_{\perp}=5\theta_{\perp}$ for the ring distribution, and at $v_{\perp}\approx 4.95\theta_{\perp}$ for the ring Maxwellian. Also, in Figure 4(f), the two shell distributions are similar around the peaks for $V=5\theta$. The maximums are located at $v=5\theta$ for the shell distribution, and at $v\approx 4.9\theta$ for the shell Maxwellian. These results suggest that the ring and shell Maxwellians are good alternatives for the ring and shell distributions with Gaussian width. On the other hand, when $V\lesssim\theta$, the ring (shell) distribution and the ring (shell) Maxwellian are very different, as shown by the dotted lines. As already stated, the artificial boundary of the ring and shell distributions becomes problematic.
@@ VIII Other isotropic distributions
We present numerical procedures for two isotropic distributions. First, we discuss the super-Gaussian distribution, also known as the self-similar distribution or the multivariate exponential power distribution. The distribution has long been discussed in laser plasmas [11, 28] but also useful in space plasmas [48]. It is given by
@@
Finally we consider the filled-shell distribution, which plays a role in a distant heliosphere. The phase space density follows a power-law $\propto v^{p}$ inside a spherical shell ($v\leq V$),
@@
The spectral index varies from problem to problem. The so-called Vasyliunas–Siscoe model [47] employs p=3/2p=-3/2 [e.g., 55], but we find other choices in the literature [e.g., $p=1$; 14].
@@
Figure 5: The phase space densities (black lines) and Monte Carlo results of $10^{6}$ particles (blue histograms) for (a) the super Gaussian distribution ($p=3$, $\theta=1$) and (b) the filled-shell distribution (p=3/2p=-3/2, $V=2$) are presented.
IX Numerical test
Figure 6: The Kullback–Leibler divergence of the numerical distributions from the analytical solutions (Eq. (108)).
Our algorithms are tested in each sections (Sections IIIVIII), either in 1-D histograms and in the 2-D colormaps, but we present one more test to verify them. We have conducted a Kullback–Leibler (KL) divergence test to our datasets, as done in our earlier study [57]. For the $i$-th bin in 1-D histograms or the $i$-th cell in 2-D colormaps, we compare exact densities $P(i)$ and normalized particle count numbers $Q(i)$. Similarity between $P$ and $Q$ is evaluated by the KL divergence, also known as the cross entropy,
$D_{\rm{KL}}(P\|Q)\equiv\sum_{i}(P(i)+\epsilon)\log\frac{P(i)+\epsilon}{Q(i)+\epsilon}$ (108)
where a small number $\epsilon=10^{-10}$ is added to avoid zero. It is non-negative, and it becomes zero when and only when the two distributions are identical.
Fig. 6 shows the KL divergence for the nine distributions, as a function of the total particle number. The parameters are detailed in the relevant sections. For the subtracted Kappa distribution, $\Delta=0.2$ is set (Fig. 3(b)). For the ring and shell distributions (Figs. 4(a,b)), the piecewise rejection method is used. Note that it is not meaningful to compare different distributions quantitatively, because the data come from sources of difference sizes: PSDs, omni-directional velocity distributions ($4\pi v^{2}f(v)$), or a 2-D colormap. Importantly, $D_{\rm{KL}}$ approaches zero in all the cases, as we increase the particle number. This suggests that the Monte Carlo results approximate the exact distributions very well.
X Discussion and Summary
@@
In our earlier papers [56, 57], we have proposed numerical procedures using three elemental variates: the uniform variate, the normal variate, and the gamma variate. In this paper, we have successfully extended this approach to nine different VDFs: the $(r,q)$ distribution, the regularized Kappa distribution, the subtracted Kappa distribution, the ring distribution, the shell distribution, the ring Maxwellian, the shell Maxwellian, the super-Gaussian, and the filled-shell distributions. With help from the basic statistical distributions (Section II), mathematics behind the recipes are fully explained. The numerical recipes for all the VDFs are provided in Tables, making it easy for users to implement them in the code. The algorithms are tested in each sections, as well as in Section IX.
@@
In Section III, we have discussed the $(r,q)$ distribution that generalizes the Kappa and flattop distributions. One can use either of the beta-prime method and the piecewise rejection method, if one is aware of limitations of each. Sections IV and V have discussed the two advanced Kappa distributions. For the regularized Kappa distribution, the two rejection methods and their preferred parameter spaces are described. The simple procedure was proposed for a very recent model of the subtracted Kappa distribution. The loss-cone filling parameter $\Delta$ is considered in Eq. (68). A previous procedure for the subtracted Maxwellian [57, Section 2] can be similarly modified to incorporate the filling parameter. In addition, after submitting this paper, Min et al. [26] independently developed the same procedure for their PIC simulation (see Appendix in their paper). This work provides a more detailed explanation to this cutting-edge method.
@@
In Section VII, we have introduced the ring Maxwellian and the shell Maxwellian. We would propose them as alternatives for the ring and shell distributions in Section VI. When the ring/shell velocity substantially exceeds the thermal width, $V\gg\theta$, these distributions are similar to the conventional ring and shell distributions. If we evaluate their differences using Eq. (100) at $V\pm\sigma$, where $\sigma=\theta/\sqrt{2}$ is the standard deviation, the two ring distributions are in agreement within 10% for $V\gtrsim 3.9\theta_{\perp}$, and the two shell distributions are in agreement within 15% for $V\gtrsim 5.25\theta$. Physically, the new Maxwellians correspond to the radial or spherical scattering of a seed population, whereas the ring and shell distributions lack a justification for their Gaussian width. When $V\lesssim\theta$, the conventional distributions become somewhat artificial near $v=0$, and probably it does not make sense to employ them. In contrast, the (ring and shell) Maxwellians have no artificial problem at $v=0$, but they are no longer ring- or shell-shaped when $V\leq\theta_{\perp}$ (ring) or $\leq\sqrt{3/2}\theta$. Practically, the new Maxwellians are easier to use. As shown in Sections VI and VII, it is much easier to estimate pressures, pressure anisotropies, and energy density for the new Maxwellians than those for the conventional distributions. Also, Monte Carlo procedures for the new Maxwellians are much simpler than those for the conventional distributions, as shown in Tables 4 and 5. These facts suggest that the ring or shell Maxwellians are favorable for modeling of heliospheric kinetic processes in ring or shell-shaped VDFs.
Technically, the proposed algorithms raise minor concerns when deployed on graphics processing units (GPUs). In particular, algorithms that rely on the rejection method can be inefficient on GPUs. This is because the number of iterations required by the rejection loop may vary from thread to thread, which is not well suited to the GPU’s execution model. Furthermore, at the time of writing, neither NVIDIA’s cuRAND library nor AMD’s hipRAND library provides gamma generators, although both libraries support uniform and normal variates. Consequently, we need to generate gamma variates on the CPU or to implement custom gamma generators on the GPU. Since GPU computation has only recently become popular in heliophysics, further investigation is needed to see whether these issues are critical and to find other potential problems.
@@
Together with our earlier works [56, 57], we can now generate a wide variety of single-component VDFs in heliophysics for particle simulations. In more complex senarios, VDFs may have multiple components, some of which are not approximated by well-known VDFs. For example, solar-wind electrons often consists of core, halo, and strahl populations. In such cases, it may be necessary not only to develop separate procedures for each component, but also to use a fully numerical inversion method [4]. This should be evaluated on a case-by-case basis.
@@
In summary, we have provided numerical procedures for generating nine different VDFs in heliophysics research. We hope that this work will make kinetic modeling of heliospheric plasma processes easier than before.
@@ Acknowledgements
Data Availability Statement
The Jupyter notebook for this article is archived in Zenodo [61].
@@ Appendix A Piecewise rejection method for log-concave distributions
The distribution has a maximum at the most probable speed:
@@
are the two switching points. As an example, the envelope functions are drawn by the dashed orange lines in Figures 4(a,b). The ratio of the areas below the envelopes is $\lambda(v_{L}):(x_{R}-x_{L}):\lambda(v_{R})$. Then a rejection algorithm can be constructed, as shown in Table 7. The acceptance efficiency is
${\rm eff}=\frac{1}{g(v_{m})\big\{(x_{R}-x_{L})+\lambda(v_{R})+\lambda(v_{L})\big\}}.$ (122)
This gives $\gtrsim 86\%$ for $V\gtrsim\theta$ and $88\%$$89\%$ for $V\gtrsim 2\theta$ for both ring and shell distributions. In practice, the normalization constant in $g(x)$ can be dropped in the procedure (Table 7).
@@
$p_{L}\leftarrow\dfrac{\lambda_{L}}{S}$$p_{R}\leftarrow\dfrac{\lambda_{R}}{S}$$p_{C}\leftarrow 1-p_{L}-p_{R}$